OpenStudy (anonymous):

Given (C^T -2I)^-1 =|2 1 | find C |-1 0 |

OpenStudy (anonymous):

how do i un-inverse?

OpenStudy (gorv):

what is T here ??

OpenStudy (anonymous):

The transpose of the matrix

OpenStudy (gorv):

ohhh....

OpenStudy (anonymous):

the 2I is 2 times the unit matrix and the ^-1 is the inverse of the entire matrix

OpenStudy (gorv):

u know how we find inverse of matrix of order 2X2

OpenStudy (gorv):

take inverse on both side

OpenStudy (dumbcow):

find inverse \[(A^{-1})^{-1} = A\]

OpenStudy (gorv):

\[(A^{-1})^{-1}=A\]

OpenStudy (anonymous):

1/|C| | b -d | |-c a |

OpenStudy (anonymous):

so the inverse on the left is |0 1| |-1 2|

OpenStudy (anonymous):

and then i just add 2I on the right and transpose it?

OpenStudy (gorv):

we also multiply by -1

OpenStudy (gorv):

and divide by determinent

OpenStudy (anonymous):

its 1/1 since |C|=2(0)-1(-1)=1

OpenStudy (gorv):

\[\left[\begin{matrix}0 & -1 \\ 1 & 2\end{matrix}\right]\]

OpenStudy (gorv):

\[C^T-2I\]

OpenStudy (gorv):

above matrix equal to this

OpenStudy (gorv):

\[2I=\left[\begin{matrix}2 & 0 \\ 0 & 2\end{matrix}\right]\]

OpenStudy (anonymous):

but if i use the formula |a b|=|b -c| |c d| |-d a|

OpenStudy (gorv):

but u forgot to multiply by -1

OpenStudy (anonymous):

then the -1 in c will swap with the 1 in b and then multiply both by -1

OpenStudy (anonymous):

i see i confused myself with the signs

OpenStudy (gorv):

we dont swap b and c

OpenStudy (anonymous):

youre right

OpenStudy (gorv):

only a and d are swapped

OpenStudy (anonymous):

i see, my formula was wrong

OpenStudy (gorv):

now after that add 2I on both side in \[C^T-2I\]

OpenStudy (gorv):

and matrix on left side

OpenStudy (anonymous):

thanks

OpenStudy (gorv):

welcome :)